OpenStudy (anonymous):

its about pressure topic... i know sound easy but... whats up with this question... but 1st i know P=F/A or maybe its just me .. lol "a 79 kg person sits on a 3.7kg chair. each leg of the chair makes contact with the floor in a circle that is 1.3cm in diameter. find the pressure exerted on the floor by eah leg of the chair, assuming the weight is evenly distributed" i've try it.. but i don't think my answer is right.. the answer for this question is 1.5 mPa

OpenStudy (anonymous):

82.7kg, thus 82.7g Newtons Divide by 4*pi*(0.013)^2

OpenStudy (anonymous):

P = F/A, F is the total force (= Weight of the person+ weight of the chair), and A is the total area (= 4* area of each leg). A point to remember is that the masses are given in Kg, Force = Weight = m*g. So, total force F is \[F =m_1\times g+m_2\times g = (m_1+m_2)\times g\]

OpenStudy (anonymous):

Also remember to divide by 100 to convert into metres BEFORE squaring, or if after divide by 100^2.

OpenStudy (anonymous):

am i finding the area of the leg using equation of \[4*\pi*d ^{2}\]

OpenStudy (anonymous):

or may \[2\pi r^{2}\]

OpenStudy (anonymous):

\[\pi r^2=\pi (d/2)^2= \pi d^2/4=Area\]

OpenStudy (anonymous):



OpenStudy (anonymous):

its really far from the real answer.. am i doing it wrong??

OpenStudy (anonymous):

To change it into m^2, did you divide by 100^2 (NOT 100)?

OpenStudy (anonymous):

\[F \approx 800N\] \[A = 4*\pi(0.013)^2 \approx 0.002\]\[P \approx 400000\]

OpenStudy (anonymous):

Yes, your area was too high

OpenStudy (anonymous):